# HD2-SSC: High-Dimension High-Density Semantic Scene Completion for Autonomous Driving

### v1

@@ Introduction
Accurate 3D geometry perception of the surrounding environment is essential for autonomous driving systems, enabling downstream tasks such as planning, navigation, and interaction with dynamic environments. 3D Semantic Scene Completion (SSC) (Roldao, De Charette, and Verroust-Blondet 2022) aims to infer both the occupancy and semantics of the 3D space, providing a voxelized understanding of the scene. Considering the complexity of outdoor environments, LiDAR-based methods (Roldao, de Charette, and Verroust-Blondet 2020; Cheng et al. 2021a; Xia et al. 2023; Li et al. 2023a; Lee et al. 2024; Wang et al. 2025) have been widely explored due to their ability to directly capture 3D structural information from input point clouds. However, they face the inherent challenges of high sensor cost and limited scalability for LiDAR data, restricting their applications to large-scale deployment in real-world scenarios.
Recently, camera-based semantic scene completion has garnered increasing attention as a cost-effective and scalable alternative utilizing 2D images as input. The pioneering work MonoScene (Cao and de Charette 2022) lifts 2D image features into dense 3D volumes with depth projection. Building on this, bird’s-eye-view (BEV) (Huang et al. 2021; Li et al. 2023b) and tri-perspective view (TVP) (Huang et al. 2023) representations are designed for improved computation efficiency. With these advances, transformer-based methods (Li et al. 2022; Wei et al. 2023) utilized voxel queries to aggregate image features into 3D scenes. VoxFormer (Li et al. 2023e) introduces a two-stage framework, aggregating 3D features of visible areas and then diffusing to the whole scene. Efforts have been made for improved performance through neighboring object structures and view spatial differences (Xue et al. 2024), context-dependent queries (Yu et al. 2024), and dense-sparse-dense design for hybrid guidance (Mei et al. 2024).
@@
LiDAR-based SSC Methods.
Due to the complexity of outdoor scenes and rich 3D structural information embedded in point clouds, LiDAR-based methods (Liong et al. 2020; Cheng et al. 2021b; Zhang et al. 2023; Ye et al. 2023) have long been the predominant solutions for SSC. UDNet (Zou et al. 2021) pioneered the use of a 3D U-Net architecture to generate scene predictions directly from LiDAR data. SGCNet (Zhang et al. 2018) enhances efficiency through spatial group convolutions for sparse 3D processing. Further advancements have focused on refining 3D scene representations, such as multi-view fusion for completing sparse scenes (Cheng et al. 2021a) and local implicit functions for continuous scene modeling (Rist et al. 2021). SCPNet (Xia et al. 2023) transfers multi-frame information to single-frame models for lightweight yet effective architectures. SSA-SC (Yang et al. 2021) and SSC-RS (Mei et al. 2023) adopt multi-branch designs to hierarchically fuse semantic and geometric features. Despite these advancements, LiDAR-based methods remain computationally expensive due to large volume of LiDAR points, limiting their scalability to real-world, large-scale applications.
@@
In recent years, camera-based SSC methods (Miao et al. 2023; Yao et al. 2023; Tong et al. 2023; Wang and Tong 2024; Wang et al. 2024; Yang and Peng 2025b, a) have gained increasing attention due to their cost-effectiveness and ease of deployment. One of the first fully visual SSC frameworks, MonoScene (Cao and de Charette 2022), pioneers a voxel-based approach by sampling image features along lines of sight. Another line of research leverages Bird’s-Eye-View (BEV) representations to aggregate multi-view image features for 3D scene modeling. LSS (Philion and Fidler 2020) introduces a frustum-based lifting mechanism, which is further enhanced by BEVDet (Huang et al. 2021), aligning multi-frame features for improved scene comprehension. BEVDepth (Li et al. 2023c) and BEVStereo (Li et al. 2023b) incorporate camera-aware depth estimation and dynamic temporal stereo information for more precise depth learning. Beyond BEV, TPVFormer (Huang et al. 2023) proposes a Tri-Perspective View (TPV) representation, decomposing voxel grids onto three orthogonal planes for efficient scene encoding. Transformer-based methods utilize BEV queries (Jiang et al. 2023) or voxel queries (Ma et al. 2024). Voxformer (Li et al. 2023e) adopts a two-stage framework with class-agnostic query proposals. Symphonize (Jiang et al. 2024) introduces a scene-from-instance paradigm to enhance feature interactions. HASSC (Wang et al. 2024) exploits the local and global hardness of each voxel, together with a self-distillation strategy for stable and consistent training process. SGN (Mei et al. 2024) presents a one-stage SSC framework with a dense-sparse-dense strategy, exploiting hybrid guidance to improve segmentation boundaries for more precise 3D semantic scene completion.
@@
Consistent with previous methods (Mei et al. 2024), we employ the ResNet-50 (He et al. 2016) network with FPN (Lin et al. 2017) to extract 2D features $F^{\rm 2D}\in\mathbb{R}^{N_{t}\times C\times H\times W}$ from input RGB images, where $N_{t}$ is the image number of temporal inputs, $C$ is the feature channel and $(H,W)$ denotes the image resolution.
@@
Following (Mei et al. 2024), we construct 3D features by sampling 2D features via 3D-2D projection mapping with camera parameters. Let $x\in\mathbb{R}^{H\times W\times Z\times 3}$ denote the centroid of $H\times W\times Z$ voxels in world coordinates. We establish the projection mapping $\pi(x)$ from pixel $(u,v)$ using the camera intrinsic parameter $K$ and extrinsic parameter $T=[R,t]$:
@@
Table 1: Camera-based 3D semantic scene completion results on the SemanticKITTI (Behley et al. 2019) validation set. denotes the methods employing EfficientNet-B7 (Tan and Le 2019) as image backbone, and represents the methods utilizing ResNet-50 (He et al. 2016) as image backbone. Best results are highlights in bold, and second-best scores are underlined.
@@
MonoScene (Cao and de Charette 2022) 36.86 11.08 23.26 0.61 0.45 6.98 1.48 1.86 1.20 0.00 56.52 14.27 26.72 0.46 14.09 5.84 17.89 2.81 29.64 4.14 2.25
TPVFormer (Huang et al. 2023) 35.61 11.36 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52
OccFormer (Zhang, Zhu, and Du 2023) 36.50 13.46 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86
IAMSSC (Xiao et al. 2024) 44.29 12.45 26.26 0.60 0.15 8.74 5.06 1.32 3.46 0.01 54.55 16.02 25.85 0.70 17.38 6.86 24.63 4.95 30.13 6.35 3.56
@@
VoxFormer (Li et al. 2023e) 44.15 13.35 26.54 1.28 0.56 7.26 7.81 1.93 1.97 0.00 53.57 19.69 26.52 0.42 19.54 7.31 26.10 6.10 33.06 9.15 4.94
DepthSSC (Yao and Zhang 2023) 45.84 13.28 25.94 0.35 1.16 6.02 7.50 2.58 6.32 0.00 55.38 18.76 27.04 0.92 19.23 8.46 26.37 4.52 30.19 7.42 4.09
Symphonies (Jiang et al. 2024) 41.92 14.89 28.68 2.54 2.82 20.44 13.89 3.52 2.24 0.00 56.37 15.28 27.58 0.95 21.64 8.40 25.72 6.60 30.87 9.57 5.76
HASSC (Wang et al. 2024) 44.58 14.74 27.33 1.07 1.14 17.06 8.83 2.25 4.09 0.00 57.23 19.89 29.08 1.26 20.19 7.95 27.01 7.71 33.95 9.20 4.81
H2GFormer (Wang and Tong 2024) 44.69 14.29 28.21 0.95 0.91 6.80 9.32 1.15 0.10 0.00 57.00 21.74 29.37 0.34 20.51 7.98 27.44 7.80 36.26 9.88 5.81
CGFormer (Yu et al. 2024) 45.99 16.87 34.32 4.61 2.71 19.44 7.67 2.38 4.08 0.00 65.51 20.82 32.31 0.16 23.52 9.20 26.93 8.83 39.54 10.67 7.84
SGN (Mei et al. 2024) 46.21 15.32 33.31 0.61 0.46 6.03 9.84 0.47 0.10 0.00 59.10 19.05 29.41 0.33 25.17 9.96 28.93 9.58 38.12 13.25 7.32
@@
Table 2: Camera-based 3D semantic scene completion results on the SSCBench-KITTI-360 (Behley et al. 2019) test set. denotes the methods employing EfficientNet-B7 (Tan and Le 2019) as image backbone, and represents the methods utilizing ResNet-50 (He et al. 2016) as image backbone. Best results are highlights in bold, and second-best scores are underlined.
@@
MonoScene (Cao and de Charette 2022) 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
GaussianFormer2 (Huang et al. 2024) 38.31 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
TPVFormer (Huang et al. 2023) 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer (Zhang, Zhu, and Du 2023) 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
IAMSSC (Xiao et al. 2024) 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
@@
VoxFormer (Li et al. 2023e) 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
DepthSSC (Yao and Zhang 2023) 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.77 5.24 3.51
Symphonies (Jiang et al. 2024) 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (Yu et al. 2024) 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
SGN (Mei et al. 2024) 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86
@@
In accordance with existing camera-based SSC methods (Mei et al. 2024; Li et al. 2023e), extensive experiments and analyses are conducted to validate our HD2-SSC framework on the popular SemanticKITTI (Behley et al. 2019) and SSCBench-KITTI-360 (Li et al. 2023d) datasets.
@@
(1) The SemanticKITTI (Behley et al. 2019) dataset, based on the KITTI Odometry Benchmark (Geiger, Lenz, and Urtasun 2012), contains 22 autonomous driving sequences with 20 semantic classes and annotations. We follow the official protocol to split the total sequences into (00-07, 09-10) / (08) / (11-21) for the training / validation / test sets respectively. (2) The SSCBench-KITTI-360 (Li et al. 2023d) dataset consists of 9 densely annotated sequences of urban driving scenes. It is separated into a training set from sequences (00, 02-05, 07, 10), a validation set from sequence (06), and a test set from sequence (09). For both datasets, the semantic labels are within the range of $0\sim 51.2m,-25.6\sim 25.6m,-2\sim 4.4m]$, and the target voxel grids are represented as $256\times 256\times 32$ voxel grids with the resolution of 0.2mm.
@@
Following (Mei et al. 2024), we adopt the Intersection over Union (IoU) of occupied voxels as the evaluation metric for the class-agnostic scene completion (SC) task, and the mean Intersection over Union (mIoU) metric for the semantic scene completion (SSC) task.
@@
We crop the input RGB images of cam2 to size $1220\times 370$ for SemanticKITTI and images of cam1 to size $1408\times 376$ for SSCBench-KITTI-360, extracting 2D feature maps with $1/16$ of input resolutions for subsequent process. The feature dimension $C$ is set to 128. The size $H\times W\times Z$ of the 3D feature volume is $128\times 128\times 16$, and the final predictions are upsampled to $256\times 256\times 32$. The channel dimensions for 2D pixel features and 3D voxel features are set to $C_{\rm 2D}=256$ and $C_{\rm 3D}=32$, respectively. The expanded dimension in the pseudo voxelization block is $D_{\rm exp}=4$, and the number of pixel and voxel query vectors is $N_{\rm query}=100$. The number of geometric and semantic critical voxels is set as $k=4096$. The above hyperparameters are practically selected by tuning on the validation set. HD2-SSC is trained for 24 epochs on 4 A6000 GPUs with a total batchsize of 4. The AdamW (Loshchilov 2017) optimizer is utilized with an initial learning rate of 2e-4 and a weight decay of 1e-2.
@@
To further assess the effectiveness of HD2-SSC, we conduct an ablation study on the SemanticKITTI validation set, evaluating the impact of key components. We incrementally integrate the HSD and HOR modules into the baseline method (Li et al. 2023e), with results summarized in Table 3. Incorporating HSD yields improvements of 2.30% in IoU and 2.23% in mIoU over the baseline. Adding the HOR module enhances SSC performance over the baseline by 1.92% in IoU and 2.77% in mIoU. These results underscore the benefits of decoupling coarse pixel semantics and refining contextual details with the aligned geometric and semantic distribution. Notably, HSD and HOR modules complement each other, with their combined application achieving superior results.
@@
Figure 5 provides the visualization results from the SemanticKITTI validation set. To better illustrate the effectiveness of our HD2-SSC approach, we also present the visualization results of the best comparison method SGN (Mei et al. 2024) and the corresponding ground truth illustrations. We outline the occupancy ground truth with blue boxes in the second row as a reference for comparison. The red boxes in the fourth row indicate the area of false voxel predictions where the SGN struggles, while the green boxes in the third row highlight the improved scene completion results from our HD2-SSC approach with completion of missing voxels and correction of erroneous ones. This improvement is attributed to HD2-SSC addressing the dimension and density gap with decoupled pixel semantics and refined voxel predictions.

### v2

@@ Introduction
Accurate 3D geometry perception of the surrounding environment is essential for autonomous driving systems, enabling downstream tasks such as planning, navigation, and interaction with dynamic environments. 3D Semantic Scene Completion (SSC) (roldao20223d) aims to infer both the occupancy and semantics of the 3D space, providing a voxelized understanding of the scene. Considering the complexity of outdoor environments, LiDAR-based methods (roldao2020lmscnet; cheng2021s3cnet; xia2023scpnet; li2023lode; lee2024semcity; wang2025l2cocc) have been widely explored due to their ability to directly capture 3D structural information from input point clouds. However, they face the inherent challenges of high sensor cost and limited scalability for LiDAR data, restricting their applications to large-scale deployment in real-world scenarios.
Recently, camera-based semantic scene completion has garnered increasing attention as a cost-effective and scalable alternative utilizing 2D images as input. The pioneering work MonoScene (cao2022monoscene) lifts 2D image features into dense 3D volumes with depth projection. Building on this, bird’s-eye-view (BEV) (huang2021bevdet; li2023bevstereo) and tri-perspective view (TVP) (huang2023tri) representations are designed for improved computation efficiency. With these advances, transformer-based methods (li2022bevformer; wei2023surroundocc) utilized voxel queries to aggregate image features into 3D scenes. VoxFormer (li2023voxformer) introduces a two-stage framework, aggregating 3D features of visible areas and then diffusing to the whole scene. Efforts have been made for improved performance through neighboring object structures and view spatial differences (xue2024bi), context-dependent queries (yu2024context), and dense-sparse-dense design for hybrid guidance (mei2024camera).
@@
LiDAR-based SSC Methods
Due to the complexity of outdoor scenes and rich 3D structural information embedded in point clouds, LiDAR-based methods (liong2020amvnet; cheng20212; zhang2023attention; ye2023lidarmultinet) have long been the predominant solutions for SSC. UDNet (zou2021up) pioneered the use of a 3D U-Net architecture to generate scene predictions directly from LiDAR data. SGCNet (zhang2018efficient) enhanced efficiency through spatial group convolutions for sparse 3D processing. Further advancements have focused on refining 3D scene representations, such as multi-view fusion for completing sparse scenes (cheng2021s3cnet) and local implicit functions for continuous scene modeling (rist2021semantic). SCPNet (xia2023scpnet) transferred multi-frame information to single-frame models for lightweight architectures. SSA-SC (yangsemantic) and SSC-RS (mei2023ssc) adopted multi-branch designs to hierarchically fuse semantic and geometric features. Despite the advancements, LiDAR-based methods remain computationally expensive due to large volume of LiDAR points, limiting their scalability to real-world, large-scale applications.
@@
In recent years, camera-based SSC methods (miao2023occdepth; yao2023ndc; tong2023scene; wang2024h2gformer; wang2024not; yang2025sphere; yang2025gala) have gained increasing attention due to their cost-effectiveness and ease of deployment. One of the first fully visual SSC frameworks, MonoScene (cao2022monoscene), pioneers a voxel-based approach by sampling image features along lines of sight. Another line of research leverages Bird’s-Eye-View (BEV) representations to aggregate multi-view image features for 3D scene modeling. LSS (philion2020lift) introduces a frustum-based lifting mechanism, which is further enhanced by BEVDet (huang2021bevdet), aligning multi-frame features for improved scene comprehension. BEVDepth (li2023bevdepth) and BEVStereo (li2023bevstereo) incorporate camera-aware depth estimation and dynamic temporal stereo information for more precise depth learning. Beyond BEV, TPVFormer (huang2023tri) proposes a Tri-Perspective View (TPV) representation, decomposing voxel grids onto three orthogonal planes for efficient scene encoding. Transformer-based methods utilize BEV queries (jiang2023polarformer) or voxel queries (ma2024cotr). Voxformer (li2023voxformer) adopts a two-stage framework with class-agnostic query proposals. Symphonize (jiang2024symphonize) introduces a scene-from-instance paradigm to enhance feature interactions. HASSC (wang2024not) exploits the local and global hardness of each voxel, together with a self-distillation strategy for stable and consistent training process. SGN (mei2024camera) presents a one-stage SSC framework with a dense-sparse-dense strategy, exploiting hybrid guidance to improve segmentation boundaries for more precise 3D semantic scene completion.
@@
Consistent with previous methods (mei2024camera), we employ the ResNet-50 (he2016deep) network with FPN (lin2017feature) to extract 2D features $F^{\rm 2D}\in\mathbb{R}^{N_{t}\times C\times H\times W}$ from input RGB images, where $N_{t}$ is the image number of temporal inputs, $C$ is the feature channel and $(H,W)$ denotes the image resolution.
@@
Following (mei2024camera), we construct 3D features by sampling 2D features via 3D-2D projection mapping with camera parameters. Let $x\in\mathbb{R}^{H\times W\times Z\times 3}$ denote the centroid of $H\times W\times Z$ voxels in world coordinates. We establish the projection mapping $\pi(x)$ from pixel $(u,v)$ using the camera intrinsic parameter $K$ and extrinsic parameter $T=[R,t]$:
@@
MonoScene (cao2022monoscene) 36.86 11.08 23.26 0.61 0.45 6.98 1.48 1.86 1.20 0.00 56.52 14.27 26.72 0.46 14.09 5.84 17.89 2.81 29.64 4.14 2.25
TPVFormer (huang2023tri) 35.61 11.36 23.81 0.36 0.05 8.08 4.35 0.51 0.89 0.00 56.50 20.60 25.87 0.85 13.88 5.94 16.92 2.26 30.38 3.14 1.52
OccFormer (zhang2023occformer) 36.50 13.46 25.09 0.81 1.19 25.53 8.52 2.78 2.82 0.00 58.85 19.61 26.88 0.31 14.40 5.61 19.63 3.93 32.62 4.26 2.86
IAMSSC (xiao2024instance) 44.29 12.45 26.26 0.60 0.15 8.74 5.06 1.32 3.46 0.01 54.55 16.02 25.85 0.70 17.38 6.86 24.63 4.95 30.13 6.35 3.56
@@
VoxFormer (li2023voxformer) 44.15 13.35 26.54 1.28 0.56 7.26 7.81 1.93 1.97 0.00 53.57 19.69 26.52 0.42 19.54 7.31 26.10 6.10 33.06 9.15 4.94
DepthSSC (yao2023depthssc) 45.84 13.28 25.94 0.35 1.16 6.02 7.50 2.58 6.32 0.00 55.38 18.76 27.04 0.92 19.23 8.46 26.37 4.52 30.19 7.42 4.09
Symphonies (jiang2024symphonize) 41.92 14.89 28.68 2.54 2.82 20.44 13.89 3.52 2.24 0.00 56.37 15.28 27.58 0.95 21.64 8.40 25.72 6.60 30.87 9.57 5.76
HASSC (wang2024not) 44.58 14.74 27.33 1.07 1.14 17.06 8.83 2.25 4.09 0.00 57.23 19.89 29.08 1.26 20.19 7.95 27.01 7.71 33.95 9.20 4.81
H2GFormer (wang2024h2gformer) 44.69 14.29 28.21 0.95 0.91 6.80 9.32 1.15 0.10 0.00 57.00 21.74 29.37 0.34 20.51 7.98 27.44 7.80 36.26 9.88 5.81
CGFormer (yu2024context) 45.99 16.87 34.32 4.61 2.71 19.44 7.67 2.38 4.08 0.00 65.51 20.82 32.31 0.16 23.52 9.20 26.93 8.83 39.54 10.67 7.84
SGN (mei2024camera) 46.21 15.32 33.31 0.61 0.46 6.03 9.84 0.47 0.10 0.00 59.10 19.05 29.41 0.33 25.17 9.96 28.93 9.58 38.12 13.25 7.32
@@
Table 1: Camera-based 3D semantic scene completion results on the SemanticKITTI (behley2019semantickitti) validation set. denotes the methods employing EfficientNet-B7 (tan2019efficientnet) as image backbone, and represents the methods utilizing ResNet-50 (he2016deep) as image backbone. Best results are highlights in bold, and second-best scores are underlined.
@@
MonoScene (cao2022monoscene) 37.87 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
GaussianFormer2 (huang2024probabilistic) 38.31 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
TPVFormer (huang2023tri) 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer (zhang2023occformer) 40.27 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
IAMSSC (xiao2024instance) 41.80 12.97 18.53 2.45 1.76 5.12 3.92 3.09 47.55 10.56 28.35 4.12 31.53 6.28 29.17 15.24 8.29 7.01 6.35 4.19
@@
VoxFormer (li2023voxformer) 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
DepthSSC (yao2023depthssc) 40.85 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.77 5.24 3.51
Symphonies (jiang2024symphonize) 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
CGFormer (yu2024context) 48.07 20.05 29.85 3.42 3.96 17.59 6.79 6.63 63.85 17.15 40.72 5.53 42.73 8.22 38.80 24.94 16.24 17.45 10.18 6.77
SGN (mei2024camera) 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86
@@
Table 2: Camera-based 3D semantic scene completion results on the SSCBench-KITTI-360 (behley2019semantickitti) test set. denotes the methods employing EfficientNet-B7 (tan2019efficientnet) as image backbone, and represents the methods utilizing ResNet-50 (he2016deep) as image backbone. Best results are highlights in bold, and second-best scores are underlined.
@@
In accordance with existing camera-based SSC methods (mei2024camera; li2023voxformer), extensive experiments and analyses are conducted to validate our HD2-SSC framework on the popular SemanticKITTI (behley2019semantickitti) and SSCBench-KITTI-360 (li2023sscbench) datasets.
@@
(1) The SemanticKITTI (behley2019semantickitti) dataset, based on the KITTI Odometry Benchmark (geiger2012we), contains 22 autonomous driving sequences with 20 semantic classes and annotations. We follow the official protocol to split the total sequences into (00-07, 09-10) / (08) / (11-21) for the training / validation / test sets respectively. (2) The SSCBench-KITTI-360 (li2023sscbench) dataset consists of 9 densely annotated sequences of urban driving scenes. It is separated into a training set from sequences (00, 02-05, 07, 10), a validation set from sequence (06), and a test set from sequence (09). For both datasets, the semantic labels are within the range of $0\sim 51.2m,-25.6\sim 25.6m,-2\sim 4.4m]$, and the target voxel grids are represented as $256\times 256\times 32$ voxel grids with the resolution of 0.2mm.
@@
Following (mei2024camera), we adopt the Intersection over Union (IoU) of occupied voxels as the evaluation metric for the class-agnostic scene completion (SC) task, and the mean Intersection over Union (mIoU) metric for the semantic scene completion (SSC) task.
@@
We crop the input RGB images of cam2 to size $1220\times 370$ for SemanticKITTI and images of cam1 to size $1408\times 376$ for SSCBench-KITTI-360, extracting 2D feature maps with $1/16$ of input resolutions for subsequent process. The feature dimension $C$ is set to 128. The size $H\times W\times Z$ of the 3D feature volume is $128\times 128\times 16$, and the final predictions are upsampled to $256\times 256\times 32$. The channel dimensions for 2D pixel features and 3D voxel features are set to $C_{\rm 2D}=256$ and $C_{\rm 3D}=32$, respectively. The expanded dimension in the pseudo voxelization block is $D_{\rm exp}=4$, and the number of pixel and voxel query vectors is $N_{\rm query}=100$. The number of geometric and semantic critical voxels is set as $k=4096$. The above hyperparameters are practically selected by tuning on the validation set. HD2-SSC is trained for 24 epochs on 4 A6000 GPUs with a total batchsize of 4. The AdamW (loshchilov2017decoupled) optimizer is utilized with an initial learning rate of 2e-4 and a weight decay of 1e-2.
@@
To further assess the effectiveness of HD2-SSC, we conduct an ablation study on the SemanticKITTI validation set, evaluating the impact of key components. We incrementally integrate the HSD and HOR modules into the baseline method (li2023voxformer), with results summarized in Table 3. Incorporating HSD yields improvements of 2.30% in IoU and 2.23% in mIoU over the baseline. Adding the HOR module enhances SSC performance over the baseline by 1.92% in IoU and 2.77% in mIoU. These results underscore the benefits of decoupling coarse pixel semantics and refining contextual details with the aligned geometric and semantic distribution. Notably, HSD and HOR modules complement each other, with their combined application achieving superior results.
@@
Figure 5 provides the visualization results from the SemanticKITTI validation set. To better illustrate the effectiveness of our HD2-SSC approach, we also present the visualization results of the best comparison method SGN (mei2024camera) and the corresponding ground truth illustrations. We outline the occupancy ground truth with blue boxes in the second row as a reference for comparison. The red boxes in the fourth row indicate the area of false voxel predictions where the SGN struggles, while the green boxes in the third row highlight the improved scene completion results from our HD2-SSC approach with completion of missing voxels and correction of erroneous ones. This improvement is attributed to HD2-SSC addressing the dimension and density gap with decoupled pixel semantics and refined voxel predictions.